OpenStudy (anonymous):

Helppp. y=7(12/17)^x without graphing determine wether the function represents exponential growth or exponential decay. Explain how you know.

OpenStudy (campbell_st):

well look at the base of the exponent 12/17 as you raise the power... it will get smaller... so for example x = 0: (12/17)^0 = 1 x =1: (12/17)^1 = 12/17 x = 2: (12/17)^2 = 144/289 so the value decreases and the function is exponential... the base is 12/17 hope it helps

OpenStudy (anonymous):

@campbell_st Hey thanks! so basically it would be an exponential growth because 12/17>1 correct?

OpenStudy (anonymous):

No, it would be an exponential decay because 12/17<1.

OpenStudy (campbell_st):

well no... since the base is 12/17 its a decay model... \ you need to know growth models have a base > 1 or < -1 compound interest (1 + r/100)^n always bigger than 1 for decay the base is -1 < base < 1 hope it helps

OpenStudy (anonymous):

the exponent function you're studying is defined as f(x) = a(b)^x, where a ≠ 0, and b > 0 since a > 0, and 0 < b < 1, it's a decay function

OpenStudy (anonymous):

@sourwing @campbell_st @dajakesta Thank youuu guys :) i understand- i was so confused with the whole decay and exponential shenanigans. Thanks